find pressure at which the PV vs P curve acquires minima for 1 mole of oxygen gas at 0°C .    a=1.36 L² atm mol‾² , b=32 cm³ mol¯¹

Grade:11

1 Answers

Ramesh V
14 years ago

for n=1, T=273K

the eqn is : PV = Pb + RT + ab/V2 - a/V  --- (1)

differentiating for minima , d(PV)/dV = 0 whihc gives

0 = -2ab/V3 + a/V2

so V = 2b  .. (2)

hence on substituting V,a,b,R,T, and putting (2) in (1)

we get P = RT/b - a/4b2

on substituting values you can easily get value of P

--

regards

Ramesh